The Constitution of the Student Information Processing Board of the Undergraduate Association of the Massachusetts Institute of Technology

As of April 2013

Article I Preamble The members of the Student Information Processing Board, in order to provide students with improved access to computational facilities, to serve as an advocate for the MIT Community to the Faculty and the Administration in computer-related topics, to function as an information center for members of the MIT Community interested in topics in computer science or computing, to conduct, manage, and support experiments run by students to advance the techniques of organization, planning, and administration of access to computers, and to initiate, develop, and provide computer services deemed valuable to the MIT Community, do establish this Constitution as a framework for conducting business to meet these goals. All affairs of this Board shall be conducted in light of these purposes with service to the students of the Massachusetts Institute of Technology always being held the uppermost among all considerations.

Article II Purpose To this effect, the Board shall administer such funds as MIT has allocated to this Board for these functions, subject in purpose and actual operations to the decisions of the MIT Administration. The Board shall also enter into negotiations with managers of computers affiliated with MIT research projects, to the end that the MIT Community may make use of these facilities. Finally, the Board shall investigate the possibility of using computers outside of MIT jurisdiction.

Article III Membership

Section 1 Requirements Full membership in the Board is open to all students of the Massachusetts Institute of Technology; associate membership is open to anyone. Any person may join and remain a member of the Board by assenting to its principles as described in this Constitution, by participating in the work of the Board, and by acceptance by the Board according to the procedure in Article IV.

Section 2 Associate Membership An associate member of the Board shall not vote in elections and shall not hold any position on the Executive Committee, but may participate fully in the work and activities of the Board. Consideration of a person for associate membership shall be accomplished by the procedure outlined in Article IV.

Section 3 Full Membership Only full members may vote in elections and hold positions on the Executive Committee. Full members shall be expected to participate in the work and activities of the Board. Consideration of a person for full membership shall be accomplished by the procedure outlined in Article IV.

Section 4 Membership Status An active member of the Board shall be defined as a member in good standing who has participated in some service furthering the goals of the Board and has attended at least one meeting during the preceding 30 days. Any member not conforming to these minimum criteria shall be classified as inactive, at the discretion of the Executive Committee.

Between the end of any academic vacation (including IAP and summer vacation) lasting one week or more and the end of the next meeting, any member active at the start of the vacation shall also be considered active.

Only active members may vote in elections and hold positions on the Executive Committee.

Section 5 Prospectives for Membership Any person interested in applying for membership in the Board shall undertake an introductory period during which that person, referred to as a prospective, may be familiarized with the operation of the Board. The introductory period shall last at least one month, and the prospective shall attend at least four general meetings of the Board during the introductory period. After such an introductory period has been served, that prospective may become an active member of the Board by the procedure in Article IV.

Article IV Election of Members

Section 1 Nomination of Prospectives In order to consider a prospective for membership in the Board, a motion nominating the prospective for membership shall be proposed and seconded by active full members present at a meeting of the Board, and accepted by the prospective. By accepting this nomination, a prospective shall assent to the principles of the Board as described in this Constitution and agree to participate in the work of the Board. This motion to consider the prospective shall be carried by a majority of the active full members of the Board present.

Section 2 Notification of Election of Prospectives The prospectives shall be elected to membership in the Board by the active full members of the Board. The meeting at which the membership elections for those nominated shall be held shall be scheduled no sooner than six (6) days following the meeting at which the motion to consider the prospectives for membership was entertained. Each active full member shall be sent a notification of the date of the elections and a list of the prospectives nominated for membership at least four (4) days in advance, unless the member waives such notification in writing to the Chair.

Section 3 Absentee and Proxy Voting for Prospectives Absentee voting shall be allowed for the purpose of electing the prospectives. Absentee ballots shall be valid only if presented to the Chair before voting commences. Proxy voting is not allowed.

Section 4 Quorum Quorum for the election shall be the number of votes needed to pass (as stated in Section 6 (Election)) plus two; absentee ballots shall be counted towards this quorum. If quorum is not achieved or if the prospective is not present, the election shall be rescheduled by the Chair, and all absentee ballots shall be voided (and must be resubmitted for the rescheduled election). If, on two separate occasions, quorum is achieved but the prospective is not present, the nomination of that prospective member shall be automatically withdrawn.

Section 5 Questioning of Prospectives Anyone present at the election may question the prospectives for membership; this questioning period must be closed before voting commences. The Chair may restrict questioning and limit discussion as the Chair deems appropriate.

Section 6 Election of Prospectives All voting shall be by secret ballot. If more than one prospective is under consideration, ballots not exactly marked with the name (first and last names, or just first name or last name if unique) or username of each prospective shall be wholly discarded as abstentions. Ballots not exactly marked with “yes”, “no”, or “abstain” for each prospective shall be wholly discarded as abstentions.

These conditions satisfied, new membership shall be granted by garnering the floor of one-half of the total active full membership and the floor of two-thirds of the active full members voting:

. _{max (⌊\frac{1}{2} × number of active full members⌋ , ⌊\frac{2}{3} × number of yes or no votes⌋)}

For the purpose of this calculation, abstentions shall not count as votes.

Section 7 Revocation of Membership Revocation of membership shall be accomplished by the procedure outlined in Article IV, subject to identical restrictions upon meeting schedules and approval required for the carry of the respective motions, except that the member in question need not be present. A member’s membership may be suspended, until this revocation procedure is complete, by a two-thirds vote of the Executive Committee. Should the revocation motion fail, the current Executive Committee shall not suspend the member’s membership again.

Article V Officers and Duties

Section 1 Duties of the Executive Committee The governance and all decision-making of the Board, subject to the limitations specified in this Constitution, shall be vested in a nine-member Executive Committee, which shall be composed of four officers plus five Members-at-Large.

The Executive Committee shall interpret by a majority vote the meaning of any portion of this Constitution should a question arise. Any interpretation shall be reported at the next general meeting of the Board.

The Executive Committee may appoint working committees from the general membership of the Board as it deems necessary. At least one member of the Executive Committee shall be on each one of these working committees. The Executive Committee shall be empowered to dissolve any working committee as it deems necessary, and shall have final review of all decisions of said committees.

Those members of the Board who shall have charge of a service of the Board shall be appointed in accord with guidelines set down by the Executive Committee. The Executive Committee shall be empowered to remove any such member for failure to perform the duties of that service or for other violations of this Constitution.

Section 2 Chair The Chair shall preside at all meetings of the Board and of the Executive Committee, shall serve as an official spokesperson for the Board, shall guide the long-term planning of the Board, shall ensure that the duties of the other officers are completed in their absence, and shall ensure that all elections proceed fairly and according to the proper procedure described in this Constitution.

Section 3 Vice-Chair The Vice-Chair shall serve as an advisor to the Chair and shall be in charge of the general operations and daily affairs of the Board, including delegating responsibility for the upkeep of services and resources. The Vice-Chair shall assume all responsibilities of the Chair in the Chair’s absence.

Section 4 Treasurer The Treasurer shall keep accurate records of all financial dealings of the Board, and shall serve in an advisory capacity to the Chair and the Board in matters pertaining to financial policy. The Treasurer shall preside at meetings in the absence of both the Chair and the Vice-Chair.

Section 5 Secretary The Secretary shall keep records of all meetings, shall direct all necessary correspondence and contact information, shall be charged with maintaining general office records and files, and shall keep records pertaining to the current active or inactive status of all members. The Secretary shall preside at meetings in the absence of the other officers.

Section 6 Members-at-Large In addition to the four officers, five Members-at-Large shall serve on the Executive Committee in advisory and voting capacities.

Article VI Election of the Executive Committee

Section 1 Date of Elections Elections of the Executive Committee shall be held annually at the second general meeting of the Spring academic term, at the discretion of the Executive Committee, and subject to the normal override procedure described in Article VIII.

Section 2 Notification of Election The Executive Committee shall be elected by the active full members of the Board. Each active full member shall be sent a notification of the date of the elections and a list of members eligible for election to the Executive Committee at least six (6) days in advance, unless the member waives such notification in writing to the Chair.

Section 3 Qualifications The members of the Executive Committee must be active full members in good standing. There shall be no campaigning before the elections.

Section 4 Election Order All elections shall be by secret ballot. Elections shall be held first for Chair, then for Vice-Chair, then for Treasurer, and finally for Secretary. Following these elections, elections shall be held for the five Members-at-Large.

Section 5 Description of Duties Preceding each election, the description of the duties of the position shall be read from Article V.

Section 6 Nominations Candidates for offices must be nominated and seconded by active full members present, and must accept the nomination. By accepting this nomination, a candidate shall agree to perform all the duties of the office if elected. All eligible active full members are automatically nominated as Members-at-Large, but may decline the nomination if they choose.

Section 7 Questioning of Officer Nominees Anyone present at the election may question the nominees for each office; this questioning period must be closed before ballots are distributed.

Section 8 Election of Officers Election of officers shall be by a majority (more than half) of those voting. For the purpose of this calculation, abstentions shall not be counted as votes. Valid ballots must be exactly marked with the name (first and last names, or just first name or last name if unique) or username of a single candidate; “yes” (in the event only a solitary candidate is running); “no”, “none”, or “none of the above”; or “abstain”. All valid ballots marked “abstain” and all invalid ballots shall be wholly discarded as abstentions.

If the ballots marked “no”, “none”, or “none of the above” garner a majority of the votes, that election and all elections following shall be postponed to a later time to be determined by the Executive Committee. Otherwise, if no candidate garners a majority of the votes, then the election shall proceed according to the following procedure after any candidates receiving zero votes in an election are automatically dropped.

If no candidate receives a majority of the votes, then all candidates in last place shall be dropped provided that the sum of votes cast for them is less than half of the votes cast. If the sum of votes cast for the last-place candidates is at least half of the votes cast, a runoff election shall be held to determine which of the last-place candidates to retain on the ballot. In this runoff, the first-place candidate(s) shall be retained on the ballot, and all other runoff candidates shall be dropped. Any additional candidates may be dropped only by unanimous consent of the active full members present. If, however, no candidates are dropped in this procedure, that election and all elections following shall be postponed to a later time to be determined by the Executive Committee.

In either case, a vote for the remaining candidates shall be held.

In case of a tie among all candidates, that election and all elections following shall be postponed to a later time to be determined by the Executive Committee.

Section 9 Election of Members-at-Large For the election of Members-at-Large, each active full member shall be allowed five votes, for five different candidates. Ballots not exactly marked with five votes, with each vote consisting of the name (first and last names, or just first name or last name if unique) or username of a candidate, shall be wholly discarded as abstentions.

Candidates receiving a plurality of votes shall be elected; that is, the five candidates receiving the largest number of votes shall be elected.

In case of a tie for last position, the candidates of the tie shall be chosen by a runoff election to be held as soon as possible. In this runoff election, each ballot must contain a choice for each remaining opening in the Executive Committee from among the candidates of the tie, and the final member or members of the Executive Committee shall be elected upon receipt of a plurality of votes.

Section 10 Absentee and Proxy Voting Absentee voting shall be allowed for the purpose of electing the Members-at-Large. Valid absentee ballots must contain at least the same number of votes as the total number of officers and Members-at-Large being elected (nine at the annual election), ranked in order for different eligible candidates. The first votes on an absentee ballot for eligible members shall be counted for the number of Members-at-Large needed. Absentee ballots shall be valid only if presented to the Chair before the election meeting is called to order. Proxy voting is not allowed.

Section 11 Start of Term of Office The newly elected Executive Committee shall take office at the next general meeting of the Board, unless a member contests an election in which case the current Executive Committee shall be allowed to remain in office for an extra week to evaluate any unresolved complaints.

Section 12 Approval of Elections Should a member contest the results of a particular election, that member shall notify the Executive Committee of the complaint within four (4) days after the election. Based on this complaint, the electoral process must be evaluated as soon as possible by the current Executive Committee. A particular election can only be dismissed by a two-thirds vote of the Executive Committee. If the Executive Committee dismisses a particular election, all the elections for the Executive Committee that followed are automatically dismissed, and new elections for those positions must be held as soon as possible. Grounds for dismissing an election shall be limited to failure to abide by the proper electoral procedure. The decision of the Executive Committee to approve or dismiss an election is subject to the normal override procedure described in Article VIII. Should the current Executive Committee fail to dismiss the contested election before the time specified in Section Section 11, then the newly elected Executive Committee shall take office and the complaint shall be automatically dismissed.

Section 13 Revocation of Membership in the Executive Committee Membership in the Executive Committee shall be automatically revoked by resignation or by loss of membership in the Board. Membership in the Executive Committee shall not be revoked by loss of MIT student status until an election, scheduled by the Executive Committee and subject to the normal override procedure described in Article VIII, has replaced the member.

Membership in the Executive Committee can be revoked according to the normal override procedure described in Article VIII.

Section 14 Filling Vacancies in the Executive Committee Vacancies occurring in the Executive Committee shall be filled by special election following the same procedure as described in Article VI, at a date determined by the Executive Committee, subject to the normal override procedure described in Article VIII.

Article VII Meetings

Section 1 General Meetings General meetings of the Board shall be held periodically, at a time and place arranged by the Executive Committee, upon notice to all members of the Board at least seven (7) days in advance. Any member of the Board may call for a general meeting at a specified time and place. The majority of the active full members of the Board must approve the meeting for it to be held, and this vote is not subject to the approval of the Executive Committee. All general meetings of the Board shall be open to the public unless a meeting voted closed to non-members of the Board by a majority of all active full members of the Board present. Nominations and elections of prospectives and of the Executive Committee may be entertained at general meetings. Business shall be conducted at general meetings subject to the approval of the Executive Committee.

Section 2 Special Meetings Any member of the Board may call for a special meeting at a specified time and place. The majority of the active full members of the Board must approve the meeting for it to be held, and this vote is not subject to the approval of the Executive Committee. Any business suitable for general meetings may be conducted at a special meeting, unless the meeting is designated for a particular purpose. A quorum for conducting business may be designated at the general meeting calling for the special meeting. Business conducted at special meetings is not subject to the approval of the Executive Committee.

Section 3 Meetings of the Executive Committee Meetings of the Executive Committee shall be held at the request of any member of the Executive Committee, or in accord with guidelines set down by the Executive Committee, at a time and place arranged by the Chair. Five members of the Executive Committee, including at least one officer, shall constitute a quorum for conducting business.

Meetings of the Executive Committee shall be open to all members of the MIT Community unless a meeting is voted closed to non-members of the Executive Committee by a unanimous consensus of all members of the Executive Committee present. Any member of the Executive Committee shall be free to invite any individual to attend its meetings, and shall be free to ask any such person to serve in an advisory capacity to the Executive Committee.

The discussion and business of all meetings of the Executive Committee shall be conducted through the consultation of the participants thereof, in a spirit such that each participant remains objective, and understands that his or her ideas, once expressed, become the property of the Executive Committee, rather than of his or her own self and desires. All business conducted at a meeting of the Executive Committee must be reported to the general membership of the Board by the next general meeting of the Board to remain valid.

Article VIII Review of Decisions In the case that a decision of the Executive Committee is felt detrimental to the goals of the Board, the general active membership of the Board may choose to overrule the decision in question. The procedure for overruling a decision of the Executive Committee shall be accomplished by the procedure outlined in Article IV, subject to identical restrictions upon meeting schedules and approval required for the carry of the respective motions, except that members of the Executive Committee need neither be present nor approve the overruling.

Article IX Relations with the MIT Community

Section 1 Discipline In cases of discipline, the Board may choose to withhold its services from certain individuals determined eligible for such treatment by the Executive Committee thereof. In such case, the person(s) deprived of services shall have the right of appeal before a meeting of the Executive Committee.

Section 2 Severe Cases In severe cases of discipline, the Executive Committee may request more formal action from the Administration or Faculty of MIT.

Section 3 Avenues for Appeal If a person is displeased with a guideline or a ruling of the Executive Committee, or with the execution thereof, or if he has important business to discuss, he may receive an audience at a forthcoming meeting.

Article X Amendments This Constitution may be amended by a two-thirds (2/3) vote of the Executive Committee of the Student Information Processing Board of MIT, subject to the review of the General Assembly of the Undergraduate Association of MIT.

Article XI Affiliations

Section 1 ASA Governance The Board agrees to abide by the rules and regulations of the Association of Student Activities, and its executive board. This constitution and amendments to it shall be subject to review by the ASA Executive Board to ensure that they are in accordance with the aforementioned rules and regulations.

With the exception of resources directly provided by the ASA to the Board, no provision of the previous paragraph shall be construed to provide the ASA any input into or control over the operations of the Board as regards physical or digital assets possessed by or entrusted to the Board, projects executed by the Board, or services provided by the Board internally or externally, including but not limited to: annual budget, office and machine room space, equipment under the control of the Board, and privileges (such as Athena account sponsorship) accorded the Board by IS&T. For the purposes of this section, the Board does not consider recruitment to be an asset, project, or service.

Should the Board decide that it no longer desires ASA recognition, or should the ASA approve derecognition of the Board, the previous paragraphs of this section shall be immediately nullified. Furthermore, all other provisions notwithstanding, the Board retains the right to separate from the ASA in perpetuity.

Section 2 MIT This Board shall seek advice from, and abide by, all relevant decisions of the Faculty and of the Administration of MIT.

Article XII Organization Upon ratification of this Constitution by the current members of the Board, and approval of the General Assembly of the Undergraduate Association of MIT, this Constitution, and the organization of the Board outlined herein, go into immediate effect.